What is the unsaturation number of each of these compounds? C7H16


1
Expert's answer
2021-09-26T12:15:33-0400

formula Ω = n - (m/2) + 1

n=7

m=16

unsaturation number for "C_7H_{16}" "= 7 - (\frac{16}{2}) + 1=0"